OpenStudy (anonymous):

(Lin Alg) is this Mapping between real vector spaces LINEAR or NON-LINEAR (please explain!!). L : R --> R defined as L(x) =2x+1 Thanks!

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

i think it is liear

OpenStudy (anonymous):

because for a definite x u find an unique L(x)

OpenStudy (anonymous):

A mapping between vector spaces is linear if: 1) L(x+y)=L(x)+L(y); where x and y are two vectors in the vector space 2) L(cx)= cL(c); c is a constant.

OpenStudy (anonymous):

Anwar A am i right?

OpenStudy (anonymous):

aNWAR THE GIVEN TWO CONDITIONS HOLDS HERE...

OpenStudy (anonymous):

thanks guys ya i kinda see it

OpenStudy (anonymous):

\(L(x+y)=2(x+y)+1=2x+2y+1\). But \(L(x)+L(y)=2x+1+2y+1=2x+2y+2\). Since \(L(x+y)\ne L(x)+L(y)\). The mapping in NOT linear.

OpenStudy (anonymous):

@ dipankarstudy: The first condition is not hold.

OpenStudy (anonymous):

true lol

OpenStudy (anonymous):

yes.....thanks guys.

OpenStudy (anonymous):

how do i give medals or say good answer?

OpenStudy (anonymous):

Just click on "Good Answer" on the right upper corner of each comment.

OpenStudy (anonymous):

theres no such option on mine

OpenStudy (anonymous):

nvm i see it (im new to the site) thanks again!